# PIM-malloc: A Fast and Scalable Dynamic Memory Allocator for Processing-In-Memory (PIM) Architectures

### v2

@@ Abstract.
Dynamic memory allocation is essential in modern programming but remains under-supported in current PIM devices. In this work, we first conduct a design space exploration of PIM memory allocators, examining optimal metadata placement and management strategies. Building on these insights, we propose PIM-malloc, a fast and scalable allocator for real PIM hardware, improving allocation performance by . We further enhance this design with a lightweight, per-PIM core hardware cache for dynamic allocation, achieving an additional performance gain. Finally, we demonstrate the effectiveness of PIM-malloc using a dynamic graph update workload, achieving a throughput increase.
1. Introduction
@@
Recent years have seen renewed interest in Processing-In-Memory (PIM) architectures (Draper et al., 2002; Hsieh et al., 2016a; Mutlu et al., 2020; Ahn et al., 2015b; Boroumand et al., 2019, 2017; Kwon et al., 2019; Ke et al., 2020; Gao et al., 2017; Kim et al., 2016; Alian et al., 2018; Hsieh et al., 2016b; Dai et al., 2022; Boroumand et al., 2022; Chen et al., 2022; Oliveira et al., 2022; Singh et al., 2019; Xie et al., 2021; Deng et al., 2018; Lenjani et al., 2022; Fernandez et al., 2020; Kal et al., 2023; Li et al., 2017; Hadidi et al., 2017; Kwon et al., 2021a; Stone, 1970; Cho et al., 2020; Wang et al., 2023a; Asgari et al., 2021; Jang et al., 2019; Li et al., 2024; Gao and Kozyrakis, 2016; Tian et al., 2024, 2023; Nai et al., 2018; Boroumand et al., 2018; Choi et al., 2023; Heo et al., 2024; Park et al., 2024a; Yun et al., 2024; Baek et al., 2024; Fujiki et al., 2018; Gu et al., 2025; He et al., 2025), with several commercial systems now available. Memory vendors offer domain-specific PIM solutions targeting application-specific compute primitives (Kwon et al., 2023; Kim et al., 2023b; Park et al., 2024b; Lee et al., 2021; Kwon et al., 2021b; He et al., 2020; Kang et al., 2022; Shin et al., 2018), whereas UPMEM’s solution (Devaux, 2019) provides a more flexible, general-purpose programming language that enables developers to write any parallel program for PIM execution. This high degree of programmability has spurred extensive research, broadening the reach of general-purpose PIM across diverse application domains (Gómez-Luna et al., 2021b, 2022b, a; Nider et al., 2021, 2022; Lim et al., 2023; Jonatan et al., 2024; Giannoula et al., 2022a, b; Gómez-Luna et al., 2023b, 2022a, a; Gogineni et al., 2024a; Rhyner et al., 2024a, b; Gogineni et al., 2024b; Gupta et al., 2023). Given this landscape, this paper focuses on system software support, specifically dynamic memory allocation for commodity general-purpose PIM, a capability that remains insufficiently supported in current PIM devices. We observe that the following two properties of PIM pose unique challenges in designing a high-performance PIM memory allocator:
@@
(Challenge #1: Explosion of memory address spaces to manage) Contemporary PIM systems are based on a bank-level PIM architecture, where each memory bank is paired with a dedicated PIM core (Devaux, 2019). While this design effectively enhances overall memory bandwidth, each per-bank PIM core is limited to accessing only the data stored within its respective local DRAM bank. As a result, the PIM system must manage thousands of distinct memory address spaces distributed across all PIM cores. For instance, an UPMEM-PIM system with 2,560 PIM cores requires the management of 2,560 separate memory address spaces, each necessitating its own heap space and the corresponding metadata to track allocated data, e.g., allocated memory block size and free lists. Compared to CPUs and GPUs, which have only a single memory address space to manage, such “explosion” of address spaces in PIM systems poses unique challenges, as decisions must be made regarding where to store the thousands of sets of metadata for memory allocation (e.g., all in the centralized CPU memory vs. partitioned across each PIM core’s local memory, distributed across the overall system)
@@
(Challenge #2: Wimpy PIM cores that execute the memory allocation algorithm) A typical PIM system contains a heterogeneous mix of CPU cores and PIM cores, each with its own local memory address space. CPU cores are designed using the latest performance-optimized logic process technology, allowing them to operate in the GHz frequency range with sophisticated microarchitectures that exploit instruction and data-level parallelism. In contrast, current-generation PIM cores are significantly less powerful because they are fabricated using a DRAM processing technology node, operating in the several hundreds of MHz range (Devaux, 2019; Lee et al., 2021; Kwon et al., 2021b). To compensate for the reduced performance of each PIM core, PIM systems integrate thousands of PIM cores to maximize throughput. This disparity in performance characteristics raises a critical question regarding which processor (e.g., CPU cores vs. PIM cores) is most optimal for executing the sophisticated memory allocation algorithm, as this choice impacts the speed of memory allocation.
@@
Considering these critical design challenges, our first important contribution is a thorough examination and characterization of the design space of dynamic memory allocators for PIM. Based on this exploration, we design, implement, and evaluate a fast and scalable dynamic memory allocator for PIM architectures, which we henceforth refer to as PIM-malloc. Our PIM-malloc design comes in two different versions. The software-only PIM-malloc is designed to be executable on real-world, programmable PIM devices like UPMEM-PIM. Compared to a naively designed PIM memory allocator that does not consider the unique properties of PIM, our software-only PIM-malloc performs faster with high scalability. We also present our hardware/software co-designed PIM-malloc, which provides further performance benefits on top of our software-only PIM-malloc. By allocating a small, per-PIM core hardware cache, our proposed PIM-malloc function is able to achieve an additional performance improvement compared to our software-only PIM-malloc design. Overall, we believe that PIM-malloc 111PIM-malloc will be open-sourced upon the acceptance of this work. will serve as a foundation for future research, enabling PIM programmers to explore a broader spectrum of applications and unlock the full potential of general-purpose PIM.
2. Background
2.1. UPMEM-PIM: A Programmable PIM System
@@
Hardware architecture. As depicted in Figure 1, UPMEM-PIM system comprises a host CPU connected to both standard DRAM and UPMEM-PIM modules via the memory bus. Each UPMEM-PIM module adheres to the DDR4-2400 (Samsung, 2017) protocol and DIMM form factor, housing DRAM Processing Units (DPUs) per DIMM module. The DPU itself is an in-order, 14-stage pipelined, RISC-based processor clocked at MHz. These DPUs are tightly coupled with a KB scratchpad memory (WRAM), KB of instruction memory (IRAM), and a dedicated, local DRAM bank (MRAM). Leveraging fine-grained multithreading, each DPU supports the concurrent execution of up to threads (termed “tasklets” by UPMEM). While the threads can share access to a DPU’s local MRAM, WRAM, and IRAM, they are not able to address data located in other DPUs. In other words, a given DPU has its own local memory address space (i.e., each DPU’s local DRAM bank, its own WRAM and IRAM), rendering the overall PIM system to end up having distinct memory address spaces managed independently across different DPUs. This is in stark contrast to CPUs or GPUs where multiple CPU/GPU cores share a single memory address space (e.g., global memory in CUDA).
@@
Software system and programming model. UPMEM-PIM programming model is based on the single-program multiple-data (SPMD) paradigm, requiring careful data partitioning across DPUs and threads. An LLVM (Lattner and Adve, 2004; UPMEM, 2021)-based compiler toolchain facilitates the translation of UPMEM programs (written using a C-like programming language provided by UPMEM) into DPU machine code. UPMEM-PIM employs a co-processor model, reminiscent of NVIDIA’s CUDA (NVIDIA, 2020), wherein a host CPU delegates memory-intensive tasks to the DPUs integrated within the memory subsystem. This architecture necessitates the development of two distinct binaries: one for the host CPU and another for the DPUs. The host CPU is in charge of DPU allocation, program binary distribution, data partitioning and transfer, initiation of PIM program execution, and retrieval of results. The UPMEM-PIM programming model is characterized by its scratchpad-centric memory architecture. This design requires the programmer to first explicitly transfer the working set from DRAM (MRAM) to the scratchpad (WRAM). Once the data is uploaded to the scratchpad, threads can access it using load/store instructions.
@@
Dynamic memory allocation in UPMEM-PIM. UPMEM-PIM currently does not support dynamic memory allocation in its local DRAM bank (MRAM). To simplify programming, UPMEM-PIM offers a limited form of memory allocation in the 64 KB scratchpad memory (WRAM) using an API called buddy_alloc(). This function employs the buddy allocation algorithm (Knowlton, 1965) to enable dynamic allocation and deallocation of blocks with varying sizes. However, limiting memory allocation to such small 64 KB region significantly constrains programmability as well as user productivity.
2.2. Buddy Memory Allocation
@@
The buddy allocation algorithm organizes memory in power-of-2 block sizes, allowing blocks to split or merge with buddy blocks when (de)allocating memory. It helps reduce fragmentation by consolidating free blocks of the same size, preserving large contiguous memory areas. The buddy allocator manages page-level allocations in the kernel, providing a foundation for other allocators (e.g., slab allocator (Lameter, 2014)).
@@
To enable efficient division and merging of blocks, maintaining metadata is crucial in the buddy system as it indicates the allocation status of each memory block. The capacity of metadata depends on the total memory managed by the buddy allocator and the minimum allocation size. For instance, UPMEM-PIM’s buddy_alloc() manages a mere KB of heap space in scratchpad, requiring less than 512 B of metadata. As the total memory pool increases or the minimum allocation size decreases, the number of blocks and metadata entries will grow significantly. For example, applying a vanilla buddy allocation algorithm to manage a PIM core’s 32 MB of per-bank DRAM heap space incurs 512 KB of metadata for each PIM core, amounting to more than 1 GB of metadata across the 2,560 PIM cores in the system.
3. Motivation
3.1. The Need for Dynamic Memory Allocators in PIM Architectures
@@
Dynamic memory allocation is a fundamental component of modern programming, allowing flexible and efficient management of memory resources. However, current PIM programming models lack this critical functionality so employing dynamic data structures (i.e., those that can change their size at runtime, such as linked lists) remains highly challenging. Graph analytics, which is typically memory-bound and is well-suited for PIM (Zhang et al., 2018; Dai et al., 2019; Zhuo et al., 2019; Ahn et al., 2015a; Nai et al., 2017; Besta et al., 2021), serves as a compelling example. Most real-world graphs are dynamic and change its structure over time, so updating the graph data should not become a performance bottleneck. Existing CPU/GPU-based graph analytics frameworks mitigate such performance overhead by employing dynamic data structures for high-performance graph updates (Dhulipala et al., 2019; Pandey et al., 2021; Awad et al., 2020; Basak et al., 2020; Winter et al., 2018; Basak et al., 2021). Unfortunately, current PIM devices do not adequately support dynamic memory allocation and therefore cannot utilize dynamic data structures. Consequently, developers must rely on static data structures, such as compressed sparse row (CSR), which is a fixed-size format that cannot be resized at runtime. This limitation introduces significant drawbacks, which we detail below using Figure 3.
@@
(Programmability) Managing graph updates using static data structures requires cumbersome array operations. For example, adding a single edge from node 0 to node 2 (red arrow in Figure 3(a)) necessitates updating the entire NodePtr array after the first index (top of Figure 3(b)). Additionally, the EdgeIdx array must be resized and shifted. More critically, manual memory management and pointer manipulation–or, in some cases, rebuilding the entire array–place a significant burden on programmers. In contrast, a dynamic data structure would allow memory to be allocated solely for the new edges and connected via pointers (bottom of Figure 3(b)), eliminating labor-intensive updates and significantly improving programmer productivity.
(Performance) Figure 3(c) demonstrates how graph update performance changes as the size of the pre-update graph (the graph structure before the update, shown at the top of Figure 3(a)) increases (from small graphs to large graphs), while the total number of newly added graph edges remains fixed. When the pre-update graph is stored in a static CSR format, each edge insertion requires reorganizing the array, resulting in a worst-case time complexity proportional to the total size of the graph. In contrast, a dynamic data structure only requires updating the pointer variable that references the newly added data. As a result, the size of the pre-update graph has a negligible impact on graph update performance.
3.2. Design Considerations for Dynamic Memory Allocators in PIM
@@
Q1: Where should the metadata be stored? Modern PIM systems use a bank-level PIM architecture, where each DRAM bank is paired with its own dedicated PIM core. This setup significantly improves the aggregate memory bandwidth available to the PIM system but restricts each PIM core to accessing data only within its local DRAM bank. Consequently, the PIM system must manage thousands of separate memory address spaces across all PIM cores, meaning an independent set of metadata must be maintained for each address space. For instance, a UPMEM-PIM system with 2,560 PIM cores requires managing 2,560 independent address spaces, each with its own heap space and metadata for tracking allocated data. Unlike traditional CPUs and GPUs, which operate within a single address space, PIM systems present unique design challenges due to the need to manage numerous address spaces. This raises a critical question about where to store the system-wide metadata for dynamic memory allocators: either in the host CPU DRAM in a centralized manner, or distributed across the PIM cores, where each PIM core independently maintains the heap space and metadata locally within its respective DRAM bank.
@@
Q2: Which processor architecture should retain ownership of executing the memory allocation algorithm? A PIM system comprises two main computational units: the host processor and the PIM cores, which differ significantly in architecture and performance characteristics. In the UPMEM-PIM system, for instance, the CPU serves as the host processor which contains multiple high-performance cores operating at several GHz of operating frequency, with advanced microarchitectures like branch predictors and prefetchers that enable high-performance processing. In contrast, PIM cores, often numbering in the hundreds to thousands per system, are optimized for system-wide parallel computations but operate individually at much lower clock frequencies due to limitations of DRAM process technology. For example, the PIM cores in UPMEM-PIM use an in-order pipeline design and lack microarchitectural features that exploit instruction-level parallelism for high performance. This large discrepancy in processor architecture characteristics creates noticeable trade-offs and raises interesting questions about which processor architecture is most optimal for executing the dynamic memory allocation algorithm.
@@
Based on the aforementioned two key design parameters, we classify the design space of PIM memory allocators into four distinct approaches, as summarized in Table 1. For our design space exploration, we extend UPMEM-PIM’s existing scratchpad-based buddy allocator (i.e., buddy_alloc() in Section 2.2) to support (de)allocation within each PIM core’s local DRAM bank, which we henceforth refer to as the buddy_alloc_PIM_DRAM design. In buddy_alloc_PIM_DRAM, each PIM core is given a 32 MB heap space managed with a 20-level tree metadata (), allowing a minimum allocation size of 32 bytes. In the case of buddy_alloc_PIM_DRAM’s design point executed on the PIM core (“PIM-Executed”), recall from Section 2.1 that UPMEM-PIM employs a scratchpad-centric programming model, which requires the working set to be uploaded from DRAM (MRAM) to the scratchpad (WRAM) before use. Because the size of buddy_alloc_PIM_DRAM’s metadata exceeds the scratchpad capacity, our implementation of buddy_alloc_PIM_DRAM (“PIM-Executed”) maintains a software-managed metadata buffer in the scratchpad. This buffer is designed to exploit locality by caching recently accessed metadata and its neighboring entries. During the buddy allocator’s tree traversal, a miss in this software-managed buffer triggers a metadata fetch operation, transferring a contiguous block of metadata from DRAM to its buffer. We utilize buddy_alloc_PIM_DRAM as a straw-man design to underscore the critical research challenges addressed with our proposal.
@@
Figure 4 provides pseudo-code illustrating how the four design strategies in Table 1 can be utilized to implement the straw-man buddy_alloc_PIM_DRAM. For cases where the buddy allocator is executed by the PIM core (“PIM-Executed”), we implement the buddy allocation algorithm using UPMEM-PIM’s C-based programming language, allowing each PIM core to independently handle PIM memory allocation requests (pimMalloc() in Figure 4(b,d)). In such scenarios, the host CPU must explicitly request the target PIM cores to execute the pimMalloc() functions, so a separate host-side pimLaunch() function is used for “PIM-Executed” approaches (Figure 4(b,d)). As for “Host-Executed” scenarios, the CPU cores execute the buddy allocation algorithm, updating each PIM core’s metadata to reflect the heap status after servicing the PIM memory allocation request. To maximize performance, we use the pthreads library (Kerrisk, 2014) to parallelize the execution of the buddy allocator’s metadata updates (parallel for pimMalloc() in Figure 4(a,c)).
@@
Note that in the four design approaches shown in Figure 4, there are two cases where the host CPU and PIM cores must exchange important information to properly support pimMalloc()’s execution. First, in “Host-Executed” approaches, the pointer values (ptr) returned by a pimMalloc() call must be copied over to the corresponding PIM cores (HOST2PIM). This ensures that the PIM cores can utilize these pointers during their program execution (Figure 4(a,c)). Second, when metadata storage and buddy allocator execution occur in different locations (”Host-Metadata/PIM-Executed” and ”PIM-Metadata/Host-Executed”), an explicit copy of the metadata must be exchanged between the host and PIM cores (HOST2PIM in Figure 4(b) and PIM2HOST in (c)). These hostPIM data transfers are conducted using pimMemcpy(), implemented with UPMEM-PIM’s runtime data transfer API (e.g., dpu_push_xfer()).

### v3

# I Introduction

@@ Abstract
The ability to dynamically allocate memory is fundamental in modern programming languages. However, this feature is not adequately supported in current general-purpose PIM devices. To identify key design principles that PIM must consider, we conduct a design space exploration of PIM memory allocators, examining various strategies for metadata placement and management of the allocator. Based on this exploration, we introduce PIM-malloc, a fast and scalable memory allocator for general-purpose PIM that operates on real PIM hardware, achieving a improvement in memory allocation performance. This design is further enhanced with a lightweight, per-PIM core hardware cache, specifically designed for dynamic memory allocation, achieving an additional performance improvement. Finally, we demonstrate the applicability of PIM-malloc by developing several representative PIM workloads, demonstrating its effectiveness in enhancing programmability.
I Introduction
@@
Recent years have seen renewed interest in Processing-In-Memory (PIM) architectures [31, 61, 97, 2, 13, 14, 79, 70, 38, 73, 3, 62, 27, 15, 21, 108, 124, 136, 28, 88, 33, 67, 90, 56, 80, 125, 22, 132, 4, 65, 92, 37, 128, 127, 99, 12, 139, 6, 35, 112], with several commercial PIM systems now available. Memory vendors offer domain-specific PIM solutions targeting application-specific compute primitives [77, 74, 111, 87, 78, 58, 69, 122], whereas UPMEM’s solution [29] provides a more flexible, general-purpose programming language that enables developers to write any parallel program for PIM execution. This high degree of programmability has spurred extensive research, broadening the reach of general-purpose PIM across diverse application domains [52, 53, 47, 101, 100, 93, 66, 41, 42, 50, 48, 49, 45, 115, 116, 46, 55, 86, 123]. Given this landscape, this paper focuses on system software support, specifically dynamic memory allocation for commodity general-purpose PIM, a capability that remains insufficiently supported in current PIM devices. We observe that the following two properties of PIM pose unique challenges in designing a high-performance PIM memory allocator:
@@
(Challenge #1: Explosion of memory address spaces to manage) Contemporary PIM systems are based on a bank-level PIM architecture, where each memory bank is paired with a dedicated PIM core [29]. While this design effectively enhances overall memory bandwidth, each per-bank PIM core is limited to accessing only the data stored within its respective local DRAM bank. As a result, the PIM system must manage thousands of distinct memory address spaces, one for each PIM core. For instance, an UPMEM-PIM system with 2,560 PIM cores requires the management of 2,560 separate memory address spaces, each necessitating its own heap space and the corresponding metadata to track allocated data, e.g., allocated memory block size and free lists. Compared to CPUs and GPUs, which have only a single memory address space to manage, such “explosion” of address spaces in PIM systems poses unique challenges, as decisions must be made regarding where to store the thousands of sets of metadata for memory allocation (e.g., all in the centralized CPU memory vs. partitioned across each PIM core’s local memory).
@@
(Challenge #2: Wimpy PIM cores that execute the memory allocation algorithm) A typical PIM system contains a heterogeneous mix of CPU cores and PIM cores, each with its own local memory address space. CPU cores are designed using the latest performance-optimized logic process technology, allowing them to operate in the GHz frequency range with sophisticated microarchitectures. In contrast, current-generation PIM cores are significantly less powerful because they are fabricated using a DRAM processing technology node, operating in the several hundreds of MHz range [29, 87, 78]. To compensate for the reduced performance of each PIM core, PIM systems integrate thousands of PIM cores to maximize throughput. This disparity in performance characteristics raises a critical question regarding which processor (e.g., CPU cores vs. PIM cores) is optimal for executing the sophisticated memory allocation algorithm, as this choice impacts the speed of memory allocation.
@@
Considering these critical design challenges, our first important contribution is a thorough examination and characterization of the design space of dynamic memory allocators for PIM. Based on this exploration, we design, implement, and evaluate a fast and scalable dynamic memory allocator for PIM architectures, which we henceforth refer to as PIM-malloc. Our PIM-malloc design comes in two different versions. The software-only PIM-malloc is designed to be executable on real-world, programmable PIM devices like UPMEM-PIM. Compared to a naively designed PIM memory allocator that does not consider the unique properties of PIM, our software-only PIM-malloc performs faster with high scalability. We also present our hardware/software co-designed PIM-malloc, which provides further performance benefits on top of our software-only PIM-malloc. By allocating a small, per-PIM core hardware cache, our proposed PIM-malloc function is able to achieve an additional performance improvement compared to our software-only PIM-malloc design. Overall, we believe that PIM-malloc will serve as a foundation for future research, enabling PIM programmers to explore a broader spectrum of applications and unlock the full potential of general-purpose PIM111Open-sourced at https://github.com/VIA-Research/PIM-malloc..
II Background
II-A UPMEM-PIM: A Programmable PIM System
@@
Hardware architecture. UPMEM-PIM comprises a host CPU connected to both standard DRAM and PIM modules via the memory bus (Figure 1). Each PIM module adheres to the DDR4-2400 [117] protocol and DIMM form factor, housing DRAM Processing Units (DPUs). The DPU itself is an in-order, RISC-based processor clocked at MHz. These DPUs are tightly coupled with a KB scratchpad memory (WRAM), KB of instruction memory (IRAM), and a dedicated, local DRAM bank (MRAM). Leveraging fine-grained multithreading, each DPU supports the concurrent execution of up to threads. While the threads can share access to a DPU’s local MRAM, WRAM, and IRAM, they are not able to address data located in other DPUs. In other words, a given DPU has its own local memory address space (i.e., each DPU’s local DRAM bank, its own WRAM and IRAM), rendering the overall PIM system to end up having distinct memory address spaces managed independently across different DPUs. This is in stark contrast to CPUs or GPUs where multiple CPU/GPU cores share a single memory address space (e.g., global memory in CUDA).
@@
Software system and programming model. UPMEM-PIM programming model is based on the single-program multiple-data (SPMD) paradigm, requiring careful data partitioning across DPUs and threads. An LLVM [82, 130]-based compiler toolchain facilitates the translation of UPMEM programs (written using a C-like programming language provided by UPMEM) into DPU machine code. UPMEM-PIM employs a co-processor model, reminiscent of NVIDIA’s CUDA [104], wherein a host CPU delegates memory-intensive tasks to the DPUs integrated within the memory subsystem. This architecture necessitates the development of two distinct binaries: one for the host CPU and another for the DPUs. The host CPU is in charge of DPU allocation, program binary distribution, data partitioning and transfer, initiation of PIM program execution, and retrieval of results. The UPMEM-PIM programming model is characterized by its scratchpad-centric memory architecture. This design requires the programmer to first explicitly transfer the working set from DRAM (MRAM) to the scratchpad (WRAM). Once the data is uploaded to the scratchpad, threads can access it using load/store instructions.
@@
Dynamic memory allocation in UPMEM-PIM. UPMEM-PIM currently does not support dynamic memory allocation in its local DRAM bank (MRAM). To simplify programming, UPMEM-PIM offers a limited form of memory allocation in the 64 KB scratchpad memory (WRAM) using an API called buddy_alloc(). This function employs the buddy allocation algorithm [75] to enable dynamic allocation and deallocation of blocks with varying sizes. However, limiting memory allocation to such small 64 KB region significantly constrains programmability as well as user productivity.
II-B Buddy Memory Allocation
@@
The buddy allocation algorithm organizes memory in power-of-2 block sizes, allowing blocks to split or merge with buddy blocks when (de)allocating memory. It helps reduce fragmentation by consolidating free blocks of the same size, preserving large contiguous memory areas. The buddy allocator manages page-level allocations in the kernel, providing a foundation for other allocators (e.g., slab allocator [81]).
@@
The metadata overhead of the buddy allocator grows rapidly with larger heap and finer allocation granularity. For example, UPMEM‑PIM’s buddy_alloc() oversees only KB of scratchpad heap and needs less than 512 B of metadata. In contrast, applying a vanilla buddy allocation algorithm to manage a PIM core’s 32 MB of per-bank DRAM heap space incurs 512 KB of metadata for each PIM core, amounting to more than 1 GB of metadata across the 2,560 PIM cores.
III Motivation
III-A The Need for Dynamic Memory Allocators in PIM
Dynamic memory allocation is a fundamental component of modern programming, allowing flexible and efficient management of memory resources. However, current programming models for general-purpose PIM lack this functionality (i.e., memory allocation is only supported limitedly at the small scratchpad in UPMEM-PIM), which not only complicates PIM programming but can also cause excessive memory consumption. Below, we present two concrete examples that illustrate the inefficiencies that stem from this limitation.
@@ III-A The Need for Dynamic Memory Allocators in PIM
(Case study #1: Dynamic graph update) Graph analytics, which is typically memory-bound and well-suited for PIM [140, 26, 143, 1, 98, 10], serves as a compelling example. Most real-world graphs are dynamic and change their structure over time, so updating the graph data should not become a performance bottleneck. Existing CPU/GPU-based graph analytics frameworks mitigate such performance overhead by employing dynamic data structures (i.e., those that can change their size at runtime, such as linked lists) for high-performance graph updates [30, 109, 5, 7, 134, 8]. Unfortunately, current PIM devices do not adequately support dynamic memory allocation and therefore cannot utilize dynamic data structures. Consequently, developers must rely on static data structures, such as compressed sparse row (CSR), which is a fixed-size format that cannot be resized at runtime. This limitation introduces significant drawbacks, which we detail below using Figure 3.
@@
Managing graph updates using static data structures requires cumbersome array operations. For example, adding a single edge from node 0 to node 2 (red arrow in Figure 3(a)) necessitates updating the entire NodePtr array after the first index (top of Figure 3(b)). Additionally, the EdgeIdx array must be resized and shifted. More critically, manual memory management and pointer manipulation–or, in some cases, rebuilding the entire array–place a significant burden on programmers. In contrast, a dynamic data structure would allow memory to be allocated solely for the new edges and connected via pointers (bottom of Figure 3(b)), eliminating labor-intensive updates and improving programmer productivity. Figure 3(c) demonstrates how graph update performance changes as the size of the pre-update graph (the graph structure before the update, shown at the top of Figure 3(a)) increases (from small graphs to large graphs), while the total number of newly added graph edges remains fixed. When the pre-update graph is stored in a static CSR format, each edge insertion requires reorganizing the array, resulting in a worst-case time complexity that is proportional to the total size of the graph. In contrast, a dynamic data structure only requires updating the pointer variable that references the newly added data. As a result, the size of the pre-update graph has a negligible impact on graph update performance.
(Case study #2: Attention layer in LLM inference) Large language models (LLMs) primarily consist of fully connected (FC) and attention layers, and their inference proceeds in two stages. (1) Prefill: the entire input prompt is processed in one pass, producing key–value (KV) vectors for all tokens and storing them in the KV cache. (2) Decode: it then generates output tokens auto‑regressively, appending each new KV vector to the cache (Figure 4(a)). The decode stage dominates inference latency because its sequential dataflow is memory‑bandwidth‑bound, whereas prefill is highly parallel and compute‑intensive [110]. As such, state-of-the-art LLM serving frameworks [137, 76] batch requests to enhance the arithmetic intensity of the decode stage, which helps speed up FC layers but still leaves attention layers memory‑bandwidth‑limited as each request must read and write its own KV cache.
To address the memory-bandwidth limits of attention layers, prior work [110, 25, 60, 54, 59, 89, 85] has proposed offloading attention layers to PIM while executing the remaining layers on the host processor (xPU). In such xPU+PIM systems, achieving high throughput via batching depends on minimizing the allocation size of the KV cache in PIM because KV cache storage grows linearly with both sequence length and batch size; large KV caches limit the maximum batch size that can be formed. Because the output sequence length is non‑deterministic and known only at runtime, LLM serving frameworks adopt one of two strategies for managing the KV cache. First, the framework statically reserves a large contiguous memory space and then directly manages that region through a customized memory allocator (e.g., the KV block manager in vLLM [76]). Second, it can request additional memory on demand by invoking native memory allocation functions (e.g., CUDA Virtual Memory Management [103] in vAttention [114]). Because the current PIM system software lacks adequate dynamic memory allocation support, PIM programmers must either undertake the cumbersome task of implementing a customized memory allocator from scratch or statically reserve a buffer large enough to accommodate the maximum possible number of KV caches that may be required, leading to excessive memory waste [85]. Figure 4(b) compares the maximum batch size achievable when the KV cache is managed with different memory allocation techniques. Static allocation, as implemented in PAISE [85], is constrained to a smaller maximum batch size due to poor memory utilization from fragmentation. In contrast, dynamic allocation allocates KV‑cache space economically at runtime, minimizes fragmentation, and enables larger batches.
III-B Design Space of Dynamic Memory Allocators in PIM
@@
Q1: Where should the metadata be stored? Modern PIM systems use a bank-level PIM architecture, where each DRAM bank is paired with a dedicated PIM core. This setup significantly improves aggregate memory bandwidth available to the PIM system but restricts each PIM core to accessing data only within its local DRAM bank. Consequently, the PIM system must manage thousands of separate memory address spaces across all PIM cores, meaning an independent set of metadata must be maintained for each address space. For instance, UPMEM-PIM with 2,560 PIM cores requires managing 2,560 independent address spaces, each with its own heap space and metadata for tracking allocated data. Unlike traditional CPUs and GPUs which operate within a single address space, PIM systems present unique design challenges due to the need to manage numerous address spaces. This raises a critical question about where to store the system-wide metadata for dynamic memory allocators: either in the host CPU in a centralized manner, or distributed across the PIM cores, where each PIM core independently maintains the heap space and metadata locally within its respective DRAM bank.
@@
Q2: Which processor architecture should retain ownership of executing the memory allocation algorithm? PIM system comprises of two main compute units: host processor and PIM cores, which differ significantly in architecture and performance characteristics. In UPMEM-PIM, for instance, the CPU serves as the host processor which contains multiple high-performance cores operating at several GHz of operating frequency, with advanced microarchitectures like branch predictors and prefetchers that enable high-performance processing. In contrast, PIM cores, often numbering in the hundreds to thousands per system, are optimized for system-wide parallel computations but operate individually at much lower clock frequencies due to limitations of DRAM process technology. For example, the PIM cores in UPMEM-PIM use an in-order pipeline design and lack microarchitectural features that exploit instruction-level parallelism for high performance. This large discrepancy in processor architecture characteristics creates noticeable trade-offs and raises interesting questions about which processor architecture is most optimal for executing the dynamic memory allocation algorithm.
@@
Based on the aforementioned two key design parameters, we classify the design space of PIM memory allocators into four approaches summarized in Table I. For our design space exploration, we extend UPMEM-PIM’s existing scratchpad-based buddy allocator (i.e., buddy_alloc() in Section II-B) to support (de)allocation within each PIM core’s local DRAM bank, which we henceforth refer to as the buddy_alloc_PIM_DRAM design. In buddy_alloc_PIM_DRAM, each PIM core is given a 32 MB heap space managed with a 20-level tree metadata (), allowing a minimum allocation size of 32 bytes. In the case of buddy_alloc_PIM_DRAM’s design point executed on the PIM core (“PIM-Executed”), recall from Section II-A that UPMEM-PIM employs a scratchpad-centric programming model, which requires the working set to be uploaded from DRAM (MRAM) to the scratchpad (WRAM) before use. Because the size of buddy_alloc_PIM_DRAM’s metadata exceeds the scratchpad capacity, our implementation of buddy_alloc_PIM_DRAM (“PIM-Executed”) maintains a software-managed metadata buffer in the scratchpad. This buffer is designed to exploit locality by caching recently accessed metadata and its neighboring entries. During the buddy allocator’s tree traversal, a miss in this software-managed buffer triggers a metadata fetch operation, transferring a contiguous block of metadata from DRAM to its buffer. We utilize buddy_alloc_PIM_DRAM as a straw-man design to underscore the research challenges addressed by our proposal.
@@
Figure 5 provides pseudo-code illustrating how the four design strategies in Table I can be utilized to implement the straw-man buddy_alloc_PIM_DRAM. For cases where the buddy allocator is executed by the PIM core (“PIM-Executed”), we implement the buddy allocation algorithm using UPMEM-PIM’s C-based programming language, allowing each PIM core to independently handle PIM memory allocation requests (pimMalloc() in Figure 5(b,d)). In such scenarios, the host CPU must explicitly request the target PIM cores to execute the pimMalloc() functions, so a separate host-side pimLaunch() function is used for “PIM-Executed” approaches (Figure 5(b,d)). As for “Host-Executed” scenarios, the CPU cores execute the buddy allocation algorithm, updating each PIM core’s metadata to reflect the heap status after servicing the PIM memory allocation request. To maximize performance, we use the pthreads library [71] to parallelize the execution of the buddy allocator’s metadata updates (parallel for pimMalloc() in Figure 5(a,c)).
@@
Note that in the four design approaches shown in Figure 5, there are two cases where the host CPU and PIM cores must exchange important information to support pimMalloc()’s execution. First, in “Host-Executed” approaches, the pointer values (ptr) returned by a pimMalloc() call must be copied over to the corresponding PIM cores (HOST2PIM). This ensures that the PIM cores can utilize these pointers during their program execution (Figure 5(a,c)). Second, when metadata storage and buddy allocator execution occur in different locations (“Host-Metadata/PIM-Executed” and “PIM-Metadata/Host-Executed”), an explicit copy of the metadata must be exchanged between the host and PIM cores (HOST2PIM in Figure 5(b) and PIM2HOST in (c)). These hostPIM data transfers are conducted using pimMemcpy(), implemented with UPMEM-PIM’s runtime data transfer API (e.g., dpu_push_xfer()).